BRICS Notes Series, Abstracts, 2002

July 7, 2003

This document is also available as PostScript and DVI.

References

[NS-02-8] PostScript, PDF.

Peter D. Mosses, editor. Proceedings of the Fourth International Workshop on Action Seman-

tics, AS 2002, (Copenhagen, Denmark, July 21, 2002), December 2002.

vi+133 pp. Abstract: Action Semantics www.brics.dk/Projects/AS/ is a prac-

tical framework for formal semantic description of programming languages. Since its appearance ten years ago, Action Semantics has been used to describe major languages such as Pascal, SML, ANDF, and Java, and various tools for processing action semantic descriptions have been developed. The AS 2002 workshop Proceedings includes contributed papers on recent development and applications of tool support for Action Semantics, on ways of increasing the modularity of action semantic descriptions and the definition of Action Notation, on analysis of information flow and types of actions, and on test suite generation based on Abstract State Machines. It also includes an invited paper by Egon Brger with a comparative study of computation and specification models.

[NS-02-7] PostScript, PDF.

Anders Mller. Document Structure Description 2.0. December 2002. 29 pp. Abstract: This document contains the specification of the Document

Structure Description 2.0 (DSD2) schema language for XML.

[NS-02-6] PostScript, PDF, DVI.

Aske Simon Christensen and Anders Mller. JWIG User Manual. October 2002. 35 pp. Abstract: This manual describes version 1.0 of the JWIG Web ser-

vice development system.

[NS-02-5] PDF.

Patrick Cousot, Lisbeth Fajstrup, Eric Goubault, Maurice Herlihy,

Martin Rauen, and Vladimiro Sassone, editors. Preliminary Proceedings of the Workshop on Geometry and Topology

in Concurrency Theory, GETCO ’02, (Toulouse, France, October 30–31, 2002), October 2002.

vi+97. Abstract: This volume contains the preliminary proceedings of

GETCO ’02 held in Toulouse, France, on October 30–31, 2002, as a satellite workshop of Disc 2002.

[NS-02-4] PostScript, PDF. Daniel Gudbjartsson, Anna Inglfsdttir, and Augustin Kong. An BDD-Based Implementation of the Allegro Software. August 2002. 2 pp. Abstract: We have applied Multi Teminal Binary Decision Diagrams

(MTBDD’s) to parts of the algorithms sitting inside the Allegro software. The results have been really promising and we have been able to handle considerably larger pedigrees than with other methods.

[NS-02-3] PDF.

Walter Vogler and Kim G. Larsen, editors. Preliminary Proceedings of the 3rd International Workshop on Models

for Time-Critical Systems, MTCS ’02, (Brno, Czech Republic, August 24, 2002), August 2002.

vi+141 pp. Abstract: This volume contains the preliminary proceedings of

MTCS ’02 held at the University of Aalborg on August 24, 2002, as a satellite workshop of CONCUR ’02.

[NS-02-2] PDF.

Zoltn sik and Anna Inglfsdttir, editors. Preliminary Proceedings of the Workshop on Fixed Points in Com-

puter Science, FICS ’02, (Copenhagen, Denmark, July 20 and 21, 2002), June 2002.

iv+81 pp. Abstract: This volume contains the preliminary proceedings of

FICS ’01 held in Copenhagen on July 20 and 21, 2002, as a satellite workshop of LICS ’02 under FLoC ’02.

[NS-02-1] PDF.

Anders Mller and Michael I. Schwartzbach. Interactive Web Services with Java: JSP, Servlets, and JWIG. April 2002. 99 pp. Abstract: This slide collection provides an introduction to Java-

based technologies for developing interactive Web services. Part 1 describes the basic architecture of the WWW and the relation to Java. Part 2 and 3 describe Sun’s Servlet and JSP languages. Part 4 gives an overview of the JWIG high-level language developed at BRICS, and Part 5 describes the PowerForms exten.